OpenStudy (anonymous):

Help Please Will Give Medal! Use the Law of Sines to find the missing angle of the triangle Find m< B to the nearest tenth A.110.0 degrees B.153.9 degrees C.26.1 degrees D. 70.0 degrees

OpenStudy (anonymous):

triangle

OpenStudy (anonymous):

sin a/A=sin b/B

OpenStudy (anonymous):

sin 40/14 =sin c/19

OpenStudy (anonymous):

sin c = 19*sin 40/14 c= arc sin(19*sin 40/14) total angles = 180 last angle = 180-40-c

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

180-40=40

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

no....

OpenStudy (anonymous):